A production line operation is tested for filling weight accuracy using the following hypotheses.
Hypothesis Conclusion and Action
H0: μ = 16
Filling okay; keep running.
Ha: μ ≠ 16
Filling off standard; stop and adjust machine.
The sample size is 30 and the population standard deviation is
σ = 0.9.
Use
α = 0.05.
(a)
What would a type II error mean in this situation?
Accepting H0 and letting the process continue to run when actually over-filling or under-filling exists.
Accepting H0 and stopping the process when actually no over-filling or under-filling exists.
Rejecting H0 and stopping the process when actually no over-filling or under-filling exists.
Rejecting H0 and letting the process continue to run when actually over-filling or under-filling exists.
Correct: Your answer is correct.
(b)
What is the probability of making a type II error when the machine is overfilling by 0.5 ounces? (Round your answer to four decimal places.)
  

(c)
What is the power of the statistical test when the machine is overfilling by 0.5 ounces? (Round your answer to four decimal places.)
Answer #1

From the information, observe that a production line operation is tested for filling weight accuracy using the following hypo=16-(196) =16 -0.29 =15.71 Therefore, the required probability is, P(Type – Ilerror) = P(15.71 <ă <16.29| = 16.5) p(15.71-4 1